OpenStudy (anonymous):

Solve each system by elimination. I want to know how to do it, so show me each and every step so I will know how to do it for other problems ok? So thank you. 4/3 = x - 8/9y 0=24x-51-15y

OpenStudy (anonymous):

plz help.

OpenStudy (tkhunny):

Have you considered solving the first for 'x' and substituting into the second?

OpenStudy (anonymous):

4/3 = x - 8/9y 0=24x-51-15y ---------------- x-(8/9)y = 4/3 24x-15y = 51 -------------- // multiply the top equation by 9 9x-8y = 12 24x-15y = 51 -------------- // multiply the top equation by 8. multiply the bottom equation by 3. 72x -64y = 96 72x-45y = 153 --------------- subtract the bottom from the top (72-72)x + (-64-(-45))y = (96-153) 0x +(-19)y = -57 y = 3 plug back into either equation to find x.